OpenStudy (anonymous):

What is the first step needed to solve 4 over 7 the whole multiplied by x minus 5 equal negative 13? Subtract 13 from both sides Divide both sides by 7 Add 5 to both sides Multiply both sides by 4

OpenStudy (anonymous):

@ganeshie8 any ideas

OpenStudy (anonymous):

i think that it is A

OpenStudy (anonymous):

@AaronAndyson

OpenStudy (aaronandyson):

According to the given condition, \[\frac{ 4 }{ 7 }x - 5 = -13\]

OpenStudy (aaronandyson):

So ,we first add 5 to both sides. \[\frac{ 4 }{ 7 }x - 5 + 5 = -13 + 5\]

OpenStudy (aaronandyson):

Note : -5 + 5 is equal to 0

OpenStudy (aaronandyson):

Hence, \[\frac{ 4 }{ 7 }x = -13 + 5\]

OpenStudy (anonymous):

so what i had is right

OpenStudy (aaronandyson):

Wrong.

OpenStudy (aaronandyson):

It's Add 5 to both sides

OpenStudy (anonymous):

oh thanks

OpenStudy (aaronandyson):

Wait.. Let me complete my explanation ?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

what is your explanation

OpenStudy (usukidoll):

something's wrong -13+5 = -8

OpenStudy (anonymous):

oh

OpenStudy (usukidoll):

\[\frac{ 4 }{ 7 }x = -13 + 5 \] \[\frac{ 4 }{ 7 }x = -8\] and then divide 4/7 on both sides

OpenStudy (aaronandyson):

@bubbleslove1234 not at my best atm sorry....

OpenStudy (anonymous):

it is okay

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (usukidoll):

yes?

OpenStudy (aaronandyson):

-13 + 5 = 8 \[\frac{ 4 }{ 7 }x = -8 \]

OpenStudy (aaronandyson):

Multiply both sides by \[\frac{ 7 }{ 4 }\]

OpenStudy (usukidoll):

either multiply both sides by 7/4 or divide both sides by 4/7 the result should be the same

OpenStudy (usukidoll):